# UniLGL: Learning Uniform Place Recognition for FOV-limited/Panoramic LiDAR Global Localization

### v1

@@ Abstract
LiDAR-based Global Localization (LGL) is an essential ingredient for autonomous mobile robots, owing to its robustness against illumination. However, existing LGL methods typically consider only partial information (e.g., geometric features) from LiDAR observations or are designed for homogeneous LiDAR sensors, overlooking the uniformity in LGL. In this work, a uniform LGL method is proposed, termed UniLGL, which simultaneously achieves spatial and material uniformity, as well as sensor-type uniformity. The key idea of the proposed method is to encode the complete point cloud, which contains both geometric and material information, into a pair of Bird’s Eye View (BEV) images (i.e., a spatial BEV image and an intensity BEV image), thereby transforming the LGL problem into a cascaded LiDAR Place Recognition (LPR) and pose estimation problem from the perspective of image fusion. An end-to-end multi-BEV fusion network is designed to extract uniform features, equipping UniLGL with spatial and material uniformity. To ensure robust LGL across heterogeneous LiDAR sensors, a viewpoint invariance hypothesis is introduced, which replaces the conventional translation equivariance assumption commonly used in existing LPR networks and supervises UniLGL to achieve sensor-type uniformity in both global descriptors and local feature representations. Moreover, UniLGL introduces a pipeline that leverages a pre-trained single-image vision foundation model for feature extraction to enhance the multi-BEV fusion LPR network, enabling strong generalization with only a few LiDAR data for fine-tuning. Finally, based on the mapping between local features on the 2D BEV image and the point cloud, a robust global pose estimator is derived that determines the global minimum of the global pose on without requiring additional registration. To validate the effectiveness of the proposed uniform LGL, extensive benchmarks are conducted in real-world environments, and the results show that the proposed UniLGL is demonstratively competitive compared to other State-of-the-Art (SOTA) LGL methods. Furthermore, UniLGL has been deployed on diverse platforms, including full-size trucks and agile Micro Aerial Vehicles (MAVs), to enable high-precision localization and mapping as well as multi-MAV collaborative exploration in port and forest environments, demonstrating the applicability of UniLGL in industrial and field scenarios.
@@ I-A1 LPR
To overcome this limitation, Scan Context [19] is proposed, which divides the 3D space into bins in a polar coordinate and records the maximum height of the points in each bin. In this way, the 3D spatial structure of a point cloud is encoded into a compact 2D image representation. Place recognition is then achieved by measuring the similarity between the images generated from different scans. As an extension of Scan Context [19], Scan Context++ [16] proposes a novel spatial division strategy to improve lateral invariance and integrates place retrieval with coarse yaw alignment as the initial guess of point cloud registration. While many handcrafted approaches have been developed for extracting informative LPR descriptors, inspired by the remarkable success of deep networks in natural language processing and computer vision, discriminative learning-based LPR methods have emerged and demonstrated competitive performance. PointNetVLAD [14] established a foundational LPR framework by extracting local features from raw point clouds using PointNet [33] and aggregating local features into a global descriptor via a NetVLAD[12, 13] pooling layer. The resulting global descriptor is then used for LPR through K-Nearest Neighbor (KNN) search. MinkLoc3D [34] employs sparse 3D convolutions for local feature extraction and introduces a Generalized-Mean (GeM) pooling layer to aggregate the features into a global descriptor. LoGG3D-Net[17] introduces a local consistency loss to guide the LPR network in learning consistent local features across revisits, and high-order aggregation has also been explored to enhance the repeatability of global descriptors, which results in an overall improvement in LPR performance. However, these methods all rely on point cloud input, which is characterized by their sparsity and lack of orderly structure. Some methods explore the potential of image-based representations instead of direct point-based representations. OverlapNet [35] represents point clouds in Range Image View (RIV) and employs a Siamese network to estimate the overlap between two range images, enabling robust and generalizable LPR. OverlapTransformer [36] builds upon OverlapNet [35] by incorporating a transformer-based attention mechanism [37] to enhance performance. ImLPR [18] leverages the vision transformer foundation model, DINOv2 [38], to further boost LPR performance through RIV representations.
@@ I-A2 LGL
To address this challenge, a series of methods have been proposed to learn highly discriminative descriptors. LCDNet [51] presents an end-to-end LGL framework, which employs PV-RCNN [52] for robust local feature extraction and NetVLAD [12, 13] for global descriptor aggregation. LCRNet [53] introduces a novel feature extraction backbone and a pose-aware attention mechanism to jointly estimate place similarity and 6-DoF relative pose between pairs of LiDAR scans. BEV-Place[24] and BEV-Place++[15] represent point clouds using BEV images and adopt a rotation equivariant and invariant network to extract local features from the BEV images. The local features are then aggregated into a global descriptor using NetVLAD[12, 13]. BEV-Place[24] and BEV-Place++[15] demonstrate that representing point clouds with BEV images yields superior generalization capability compared to raw point clouds, particularly under viewpoint variations and scene changes. Similarly, RING# [25] proposes a feature learning architecture to simultaneously learn LPR and 3-DoF global localization, leveraging the frequency-domain BEV representation introduced in RING[49] and RING++[50].
I-A3 LRP/LGL with Uniformity
@@
Compared to spatial and material uniformity, sensor-type uniformity has received attention in only a few handcrafted descriptor-based LPR methods [26, 27, 49, 50]. As corroborated by Solid [56], designing an LPR system with sensor-type uniformity can significantly improve performance, particularly for FoV-limited LiDARs. However, most existing learning-based LPR/LGL methods [24, 15, 25, 51, 53, 17, 33, 18] follow the paradigm of position-based place recognition [28], where networks are supervised based on the geographical distance between scans to determine whether a robot revisits the same place. These approaches implicitly assume that LiDAR observations are rotation-invariant. As illustrated in Fig. 1(b), panoramic LiDARs naturally satisfy this assumption due to their omnidirectional sensing capability. In contrast, FoV-limited LiDARs’ observations at the same location under different headings may be completely disjoint. OverlapNet [35] and OverlapTransformer[36] offer an alternative by supervising the network using overlap rather than geographical distance, thus providing a more natural metric than geographical distance for point cloud similarity. However, both OverlapNet [35] and OverlapTransformer[36] represent LiDAR scans as RIV images, which limits their generalization to FoV-limited LiDARs. Specifically, to ensure that range images do not contain large regions of vacant pixels, different horizontal resolutions have to be used for panoramic and FoV-limited LiDARs, making it difficult to generalize a single model across heterogeneous LiDAR configurations.
| Capability | ISC [20] | Solid [56] | RING++ [50] | OverlapNet [35] | BEVPlace++ [15] | RING# [25] | LCDNet [51] | Logg3D-Net [17] | ImLPR [18] | UniLGL (Proposed) |
| Place Recognition | Handcrafted | Handcrafted | Handcrafted | Learning | Learning | Learning | Learning | Learning | Learning | Learning |
@@
| Point Cloud Representation | Point | Point | BEV | RIV | BEV | BEV | Point | Point | RIV | BEV |
@@ I-B Motivation and Contributions
In view of the aforementioned analysis, as summarized in Table I, most existing methods either focus exclusively on LPR[17, 18, 20] or provide only partial solutions for LGL, such as 1-DoF heading alignment[56, 35] or 3-DoF pose estimation constrained to SE(2)[15, 50, 25]. Moreover, the uniformity illustrated in Fig. 1 is commonly neglected [15, 25] or only partially considered [27, 20, 56, 35, 51, 17, 18]. These limitations motivate us to develop a uniform LGL system, called UniLGL, that enables fully 6-DoF global pose estimation over SE(3), while simultaneously preserving spatial and material uniformity as well as sensor-type uniformity. The key idea of UniLGL is to represent the 4D point cloud using two BEV images that encode both geometric and intensity information. A novel feature fusion network is then designed elaborately to learn local features and global descriptors from the BEV images, supervised under the viewpoint invariance hypothesis. Subsequently, uniform LGL is realized by combining uniform LPR using global descriptor matching with 6-DoF global pose recovery through local feature matching. To recap, the main contributions of this paper are listed as follows:
@@ II-A Learning Spatial and Material Uniformity
where and denote the -th patch of the spatial BEV image and the intensity BEV image, respectively; and represent the learnable embedding projections for the spatial and intensity BEV images. The resulting embedded patches are concatenated with the classification vector CLS into , where is the number of patches per BEV image, and is the dimension of each embedded patch. A standard learnable position embedding is employed on to retain positional information, the resulting sequence of embedding vectors serves as input to the transformer encoder.
@@
The Transformer encoder[58], composed of alternating Multi-head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) blocks, is employed to learn both the global descriptor (classification token) for LPR and the local features (local feature token) of each patch.
where is the joint-image-level classification token, and and are patch-level local feature tokens, respectively. The joint image-level classification token is treated as the unified global descriptor for LPR.
@@
(Spatial and Material Uniformity) For now, most existing LPR methods [24, 15, 25, 19, 16, 26, 27] treat LiDAR solely as a 3D geometric sensor, neglecting the material properties encoded in the intensity channel. It has been corroborated in [20, 55] that incorporating intensity information into LPR can significantly improve performance. However, although [20] and [55] incorporate both geometric and intensity cues into BEV and intensity RIV images, respectively, the projection to 2D image inevitably causes a loss of 3D geometric fidelity, as [20] retains only and [55] retains only . To mitigate the information loss, we separately preserve geometric and intensity cues using a spatial BEV image and an intensity BEV image. Built upon the single-image feature extraction backbone, ViT[57], we design a novel feature fusion network that enables joint processing of multiple images, facilitating uniform representation learning across both spatial (geometry) and material (intensity) domains.
@@ II-C LPR Meets Foundation Models
(Why LPR Meets Foundation Models) A hypothesis that has been corroborated in [64, 65, 12, 13, 66] is that the high performance of most modern place recognition can be mainly attributed to large-scale training. Over the past decade, the scale of training data used for place recognition has grown explosively, from hundreds of thousands in early works (e.g. Pittsburgh Dataset [67] powered NetVLAD[12, 13]) to millions (e.g. Google-Landmark datasets[68] powered DeLF[64] and DeLG[65]) or even tens of millions (e.g. San Francisco XL dataset powered CosPlace[66]) in recent years. However, unlike visual images that can be easily acquired at scale through map services (e.g., Google Street View), collecting large-scale point cloud data remains challenging and resource-intensive. Consequently, most existing LPR methods either rely on classical hand-crafted features[16, 27, 19, 49, 50] or lightweight models[24, 15, 25, 51] trained on relatively small-scale datasets, typically comprising only tens of thousands of samples. This motivates the introduction of foundation models pre-trained on large-scale visual data into LPR, achieving high performance through fine-tuning with only a small amount of LiDAR data.
@@ IV-B Comparison Baseline
UniLGL aims to achieve uniform global localization across spatial and material domains, as well as between FoV-limited and panoramic LiDAR observations. To demonstrate the effectiveness of the proposed method, we present detailed quantitative analyses comparing UniLGL with SOTA LPR/LGL methods, including LoGG3D-Net[17], BEVPlace++[15], and RING++[50]. Moreover, to investigate the impact of introducing spatial and material uniformity as well as sensor-type uniformity into LGL, we conduct an ablation study on the proposed UniLGL by evaluating it under various configurations. Details of the baselines are given as follows:
@@
UniLGL: the full algorithm proposed in this paper comprehensively considers spatial and material uniformity together with sensor-type uniformity.
@@ V-A Evaluation of Place Recognition
| Sequence | UniLGL | UniLGL w/o intensity | UniLGL w/o spatial | UniLGL w/o loc. feat. | BEVPlace++ | Logg3D-Net | RING++ | UniLGL Dis. Sup. | BEVPlace++ Dis. Sup. | |
| Fov-limited LiDAR | Mid_NTU_02 | 99.87/88.88 | 99.73/85.82 | 98.72/87.35 | 99.20/86.69 | 88.56/60.56 | 80.38/45.22 | 41.63/8.33 | 80.73/75.20 | 59.94/40.34 |
| Mid_NTU_10 | 95.75/86.43 | 95.82/83.87 | 95.70/83.95 | 94.60/83.49 | 94.07/60.02 | 91.59/45.04 | 60.98/9.56 | 91.78/79.44 | 92.98/64.29 | |
| Mid_NTU_13 | 98.71/88.38 | 98.58/85.39 | 98.36/87.21 | 96.71/85.42 | 93.25/64.53 | 76.37/40.62 | 42.03/6.76 | 81.24/74.14 | 56.09/44.23 | |
| Garden_01 | 97.67/83.92 | 96.81/82.61 | 97.30/83.05 | 95.15/79.77 | 95.24/59.76 | 90.67/54.18 | 86.59/15.60 | 84.33/78.46 | 84.21/35.58 | |
| Garden_02 | 98.65/84.87 | 98.52/83.12 | 98.19/82.88 | 93.87/78.09 | 94.64/63.19 | 91.02/45.84 | 86.89/16.08 | 81.59/75.70 | 81.20/37.16 | |
| Garden_03 | 98.63/85.45 | 96.15/82.87 | 97.33/77.54 | 91.57/70.55 | 91.14/56.48 | 84.92/36.64 | 72.77/15.44 | 81.57/73.76 | 82.93/35.27 | |
| Garden_04 | 99.42/86.61 | 99.18/86.35 | 98.66/83.25 | 96.14/79.77 | 99.17/59.93 | 91.05/41.76 | 81.05/19.53 | 84.88/73.34 | 87.98/32.39 | |
| Average | 98.38/86.36 | 97.82/84.29 | 97.75/83.60 | 95.32/80.54 | 93.72/60.64 | 86.57/44.19 | 67.42/13.04 | 83.73/75.72 | 80.46/41.32 | |
| Panoramic LiDAR | OS_NTU_02 | 100.0/89.93 | 100.0/86.19 | 100.0/87.57 | 100.0/92.83 | 100.0/76.56 | 98.32/74.75 | 77.72/24.52 | 98.48/94.37 | 98.33/94.09 |
| OS_NTU_10 | 99.94/88.49 | 99.75/88.95 | 100.0/88.39 | 99.92/90.64 | 99.69/76.76 | 99.10/62.91 | 78.46/26.32 | 100.0/97.23 | 100.0/97.31 | |
| OS_NTU_13 | 100.0/90.81 | 100.0/87.37 | 100.0/89.23 | 100.0/90.72 | 99.78/76.99 | 88.99/62.77 | 62.28/18.08 | 99.71/92.65 | 99.31/93.09 | |
| Snail_81R_02 | 95.62/79.79 | 95.03/76.55 | 92.73/71.82 | 96.15/88.76 | 92.08/59.95 | 72.37/30.19 | 36.09/6.03 | 98.67/89.93 | 99.50/73.61 | |
| Snail_81R_03 | 98.12/88.19 | 97.95/81.95 | 98.22/87.82 | 98.53/89.36 | 97.60/64.07 | 76.77/28.99 | 35.56/8.48 | 99.72/91.26 | 98.47/61.83 | |
| Garden_LT_01 | 98.40/87.89 | 87.30/79.01 | 87.68/82.31 | 75.35/64.21 | 74.40/63.33 | 48.96/39.26 | 86.10/39.30 | 94.76/81.20 | 76.18/35.97 | |
| Garden_LT_02 | 98.06/87.85 | 82.98/78.21 | 86.54/79.57 | 70.89/61.41 | 70.09/64.41 | 53.23/36.09 | 70.79/33.74 | 87.30/80.75 | 73.33/33.17 | |
| Garden_LT_03 | 95.08/83.31 | 94.44/79.91 | 90.19/80.34 | 84.94/77.60 | 67.55/46.29 | 53.70/43.79 | 61.88/28.28 | 89.20/81.52 | 78.73/47.56 | |
| Garden_LT_04 | 97.03/86.69 | 94.60/86.45 | 91.64/81.17 | 85.36/73.77 | 69.24/55.00 | 40.73/39.47 | 73.55/31.12 | 92.61/82.24 | 84.80/61.57 | |

### v2

@@ Abstract
LiDAR-based Global Localization (LGL) is an essential ingredient for autonomous mobile robots, owing to its robustness against illumination. However, existing LGL methods typically consider only partial information (e.g., geometric features) from LiDAR observations or are designed for homogeneous LiDAR sensors, overlooking the uniformity in LGL. In this work, a uniform LGL method is proposed, termed UniLGL, which simultaneously achieves spatial and material uniformity, as well as sensor-type uniformity. The key idea of the proposed method is to encode the complete point cloud, which contains both geometric and material information, into a pair of Bird’s Eye View (BEV) images (i.e., a spatial BEV image and an intensity BEV image), thereby transforming the LGL problem into a cascaded LiDAR Place Recognition (LPR) and pose estimation problem from the perspective of image fusion. An end-to-end multi-BEV fusion network is designed to extract uniform features, equipping UniLGL with spatial and material uniformity. To ensure robust LGL across heterogeneous LiDAR sensors, a viewpoint invariance hypothesis is introduced, which replaces the conventional translation equivariance assumption commonly used in existing LPR networks and supervises UniLGL to achieve sensor-type uniformity in both global descriptors and local feature representations. Moreover, UniLGL introduces a pipeline that leverages a pre-trained single-image vision foundation model for feature extraction to enhance the multi-BEV fusion LPR network, enabling strong generalization with only a few LiDAR data for fine-tuning. Finally, based on the mapping between local features on the 2D BEV image and the point cloud, a robust global pose estimator is derived that determines the global minimum of the global pose on without requiring additional registration. To validate the effectiveness of the proposed uniform LGL, extensive benchmarks are conducted in real-world environments, and the results show that the proposed UniLGL is demonstratively competitive compared to other State-of-the-Art (SOTA) LGL methods. Furthermore, UniLGL has been deployed on diverse platforms, including full-size trucks and agile Micro Aerial Vehicles (MAVs), to enable high-precision localization and mapping as well as multi-MAV collaborative exploration in port and forest environments, demonstrating the applicability of UniLGL in industrial and field scenarios. The code will be released at https://github.com/shenhm516/UniLGL, and a demonstration video is available at https://youtu.be/p8D-sxq8ygI.
@@ I-A1 LPR
To overcome this limitation, Scan Context [19] is proposed, which divides the 3D space into bins in a polar coordinate and records the maximum height of the points in each bin. In this way, the 3D spatial structure of a point cloud is encoded into a compact 2D image representation. Place recognition is then achieved by measuring the similarity between the images generated from different scans. As an extension of Scan Context [19], Scan Context++ [16] proposes a novel spatial division strategy to improve lateral invariance and integrates place retrieval with coarse yaw alignment as the initial guess of point cloud registration. While many handcrafted approaches have been developed for extracting informative LPR descriptors, inspired by the remarkable success of deep networks in natural language processing and computer vision, discriminative learning-based LPR methods have emerged and demonstrated competitive performance. PointNetVLAD [14] established a foundational LPR framework by extracting local features from raw point clouds using PointNet [33] and aggregating local features into a global descriptor via a NetVLAD[12, 13] pooling layer. The resulting global descriptor is then used for LPR through K-Nearest Neighbor (KNN) search. MinkLoc3D [34] employs sparse 3D convolutions for local feature extraction and introduces a Generalized-Mean (GeM) pooling layer to aggregate the local features into a global descriptor. HeLioS [18] employs a U-Net-style architecture with sparse convolution[35] to extract local features from a point cloud, which are then aggregated into LPR global descriptors through GeM pooling and the SALAD[36] attention module. LoGG3D-Net[37] introduces a local consistency loss to guide the LPR network in learning consistent local features across revisits, and high-order aggregation has also been explored to enhance the repeatability of global descriptors, which results in an overall improvement in LPR performance. However, these methods all rely on point cloud input, which is characterized by their sparsity and lack of orderly structure. Some methods explore the potential of image-based representations instead of direct point-based representations. OverlapNet [38] represents point clouds in Range Image View (RIV) and employs a Siamese network to estimate the overlap between two range images, enabling robust and generalizable LPR. OverlapTransformer [39] builds upon OverlapNet [38] by incorporating a transformer-based attention mechanism [40] to enhance performance. ImLPR [17] leverages the vision transformer foundation model, DINOv2 [41], to further boost LPR performance through RIV representations.
@@ I-A2 LGL
To address this challenge, a series of methods have been proposed to learn highly discriminative descriptors. LCDNet [54] presents an end-to-end LGL framework, which employs PV-RCNN [55] for robust local feature extraction and NetVLAD [12, 13] for global descriptor aggregation. LCRNet [56] introduces a novel feature extraction backbone and a pose-aware attention mechanism to jointly estimate place similarity and 6-DoF relative pose between pairs of LiDAR scans. SpectralGV [57] extends Logg3D-Net[37] to LGL by leveraging its local features to register point cloud pairs that are initially matched using the global descriptors produced by Logg3D-Net[37]. BEV-Place[24] and BEV-Place++[15] represent point clouds using BEV images and adopt a rotation equivariant and invariant network to extract local features from the BEV images. The local features are then aggregated into a global descriptor using NetVLAD[12, 13]. BEV-Place[24] and BEV-Place++[15] demonstrate that representing point clouds with BEV images yields superior generalization capability compared to raw point clouds, particularly under viewpoint variations and scene changes. Similarly, RING# [25] proposes a feature learning architecture to simultaneously learn LPR and 3-DoF global localization, leveraging the frequency-domain BEV representation introduced in RING[52] and RING++[53].
I-A3 LPR/LGL with Uniformity
@@
Compared to spatial and material uniformity, sensor-type uniformity has received attention in only a few handcrafted descriptor-based LPR methods [26, 27, 52, 53]. As corroborated by Solid [60], designing an LPR system with sensor-type uniformity can significantly improve performance, particularly for FoV-limited LiDARs. However, most existing learning-based LPR/LGL methods [24, 15, 25, 54, 56, 37, 33, 17] follow the paradigm of position-based place recognition [28], where networks are supervised based on the geographical distance between scans to determine whether a robot revisits the same place. These approaches implicitly assume that LiDAR observations are rotation-invariant. As illustrated in Fig. 1(b), panoramic LiDARs naturally satisfy this assumption due to their omnidirectional sensing capability. In contrast, FoV-limited LiDARs’ observations at the same location under different headings may be completely disjoint. OverlapNet [38], OverlapTransformer[39], and HeLioS[18] offer an alternative by supervising the network using overlap rather than geographical distance, thus providing a more natural metric than geographical distance for point cloud similarity. However, HeLioS[18] only considers the sensor-type uniformity at the global descriptor level by using point clouds as input to the LPR network, which neglects such uniformity at the local feature level, restricts it to LPR without metric localization. OverlapNet [38] and OverlapTransformer[39] represent LiDAR scans as RIV images, which limits their generalization to FoV-limited LiDARs. Specifically, to ensure that range images do not contain large regions of vacant pixels, different horizontal resolutions have to be used for panoramic and FoV-limited LiDARs, making it difficult to generalize a single model across heterogeneous LiDAR configurations.
Capability |
Place Recognition | Handcrafted | Handcrafted | Handcrafted | Learning | Learning | Learning | Learning | Learning | Learning | Learning | Learning | ||||||||||||||||||||||
@@
Point Cloud Representation | Point | Point | BEV | Point | RIV | RIV | BEV | BEV | Point | Point | BEV |
@@ I-B Motivation and Contributions
In view of the aforementioned analysis, as summarized in Table I, most existing methods either focus exclusively on LPR[17, 20, 18] or provide only partial solutions for LGL, such as 1-DoF heading alignment[60, 38] or 3-DoF pose estimation constrained to SE(2)[15, 53, 25]. Moreover, the uniformity illustrated in Fig. 1 is commonly neglected [15, 25] or only partially considered [27, 20, 60, 38, 54, 37, 17, 18]. These limitations motivate us to develop a uniform LGL system, called UniLGL, that enables fully 6-DoF global pose estimation over SE(3), while simultaneously preserving spatial and material uniformity as well as sensor-type uniformity. The key idea of UniLGL is to represent the 4D point cloud using two BEV images that encode both geometric and intensity information. A novel feature fusion network is then designed elaborately to learn local features and global descriptors from the BEV images, supervised under the viewpoint invariance hypothesis. Considering the strong cross-task generalization capabilities of foundation models, as illustrated by the t-SNE[61] visualization in Fig. 2, initializing the LPR network with a foundation model (without fine-tuning) endows the LPR network with an initial capability for place recognition. As shown in Table I, foundation model has rarely been explored in the context of LPR tasks. This motivates us to explore leveraging foundation models to empower the LPR network, and to bridge the domain gap between foundation models and the LPR task through fine-tuning with a small amount of LiDAR data, enabling high-performance place recognition, as illustrated in Fig. 2(c). Subsequently, uniform LGL is realized by combining uniform LPR using global descriptor matching with 6-DoF global pose recovery through local feature matching. To recap, the main contributions of this paper are listed as follows:
@@ II-A Learning Spatial and Material Uniformity
where and denote the -th patch of the spatial BEV image and the intensity BEV image, respectively; and represent the learnable embedding projections for the spatial and intensity BEV images, and is the dimension of each embedded patch. The resulting embedded patches are concatenated with the classification vector CLS into , where is the number of patches per BEV image. A standard learnable position embedding is employed on to retain positional information, the resulting sequence of embedding vectors serves as input to the transformer encoder.
@@
The Transformer encoder[63], composed of alternating Multi-head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) blocks, is employed to learn both the global descriptor (i.e., classification token) for LPR and the local features (i.e., local feature token) of each patch.
where is the joint-image-level classification token, and and are patch-level local feature tokens, respectively. The joint-image-level classification token is treated as the unified global descriptor for LPR.
@@
(Spatial and Material Uniformity) For now, most existing LPR methods [24, 15, 25, 19, 16, 26, 27, 18] treat LiDAR solely as a 3D geometric sensor, neglecting the material properties encoded in the intensity channel. It has been corroborated in [20, 59] that incorporating intensity information into LPR can significantly improve performance. However, although [20] and [59] incorporate both geometric and intensity cues into BEV and intensity RIV images, respectively, the projection to 2D image inevitably causes a loss of 3D geometric fidelity, as [20] retains only and [59] retains only . To mitigate the information loss, we separately preserve geometric and intensity cues using a spatial BEV image and an intensity BEV image. Built upon the single-image feature extraction backbone, ViT[62], we design a novel feature fusion network that enables joint processing of multiple images, facilitating uniform representation learning across both spatial (geometry) and material (intensity) domains.
@@ II-C LPR Meets Foundation Models
(Why LPR Meets Foundation Models) A hypothesis that has been corroborated in [69, 70, 12, 13, 71] is that the high performance of most modern place recognition can be mainly attributed to large-scale training. Over the past decade, the scale of training data used for place recognition has grown explosively, from hundreds of thousands in early works (e.g. Pittsburgh Dataset [72] powered NetVLAD[12, 13]) to millions (e.g. Google-Landmark datasets[73] powered DeLF[69] and DeLG[70]) or even tens of millions (e.g. San Francisco XL dataset powered CosPlace[71]) in recent years. However, unlike visual images that can be easily acquired at scale through map services (e.g., Google Street View), collecting large-scale point cloud data remains challenging and resource-intensive. Consequently, most existing LPR methods either rely on classical hand-crafted features[16, 27, 19, 52, 53] or lightweight models[24, 15, 25, 54] trained on relatively small-scale datasets, typically comprising only tens of thousands of samples. This motivates the introduction of foundation models pre-trained on large-scale visual data into LPR, achieving high performance through the use of only a small amount of LiDAR data for fine-tuning.
@@ IV-B Comparison Baseline
UniLGL aims to achieve uniform global localization across spatial and material domains, as well as between FoV-limited and panoramic LiDAR observations. To demonstrate the effectiveness of the proposed method, we present detailed quantitative analyses comparing UniLGL with SOTA LPR/LGL methods, including BEVPlace++[15], LoGG3D-Net[37], and RING++[53]. Moreover, to investigate the impact of introducing spatial and material uniformity as well as sensor-type uniformity into LGL, we conduct an ablation study on the proposed UniLGL by evaluating it under various configurations. Details of the baselines are given as follows:
@@
UniLGL: the full algorithm proposed in this paper that comprehensively considers spatial and material uniformity together with sensor-type uniformity.
@@ V-A Evaluation of Place Recognition
Sequence | UniLGL |
BEVPlace++ | Logg3D-Net | RING++ |
FoV-limited LiDAR |
Mid_NTU_02 | 99.87/88.88 | 99.73/85.82 | 98.72/87.35 | 99.20/86.69 | 88.56/60.56 | 80.38/45.22 | 41.63/8.33 | 80.73/75.20 | 59.94/40.34 | ||||||||||
Mid_NTU_10 | 95.75/86.43 | 95.82/83.87 | 95.70/83.95 | 94.60/83.49 | 94.07/60.02 | 91.59/45.04 | 60.98/9.56 | 91.78/79.44 | 92.98/64.29 | |||||||||||
Mid_NTU_13 | 98.71/88.38 | 98.58/85.39 | 98.36/87.21 | 96.71/85.42 | 93.25/64.53 | 76.37/40.62 | 42.03/6.76 | 81.24/74.14 | 56.09/44.23 | |||||||||||
Garden_01 | 97.67/83.92 | 96.81/82.61 | 97.30/83.05 | 95.15/79.77 | 95.24/59.76 | 90.67/54.18 | 86.59/15.60 | 84.33/78.46 | 84.21/35.58 | |||||||||||
Garden_02 | 98.65/84.87 | 98.52/83.12 | 98.19/82.88 | 93.87/78.09 | 94.64/63.19 | 91.02/45.84 | 86.89/16.08 | 81.59/75.70 | 81.20/37.16 | |||||||||||
Garden_03 | 98.63/85.45 | 96.15/82.87 | 97.33/77.54 | 91.57/70.55 | 91.14/56.48 | 84.92/36.64 | 72.77/15.44 | 81.57/73.76 | 82.93/35.27 | |||||||||||
Garden_04 | 99.42/86.61 | 99.18/86.35 | 98.66/83.25 | 96.14/79.77 | 99.17/59.93 | 91.05/41.76 | 81.05/19.53 | 84.88/73.34 | 87.98/32.39 | |||||||||||
Average | 98.38/86.36 | 97.82/84.29 | 97.75/83.60 | 95.32/80.54 | 93.72/60.64 | 86.57/44.19 | 67.42/13.04 | 83.73/75.72 | 80.46/41.32 | |||||||||||
Panoramic LiDAR |
OS_NTU_02 | 100.0/89.93 | 100.0/86.19 | 100.0/87.57 | 100.0/92.83 | 100.0/76.56 | 98.32/74.75 | 77.72/24.52 | 98.48/94.37 | 98.33/94.09 | ||||||||||
OS_NTU_10 | 99.94/88.49 | 99.75/88.95 | 100.0/88.39 | 99.92/90.64 | 99.69/76.76 | 99.10/62.91 | 78.46/26.32 | 100.0/97.23 | 100.0/97.31 | |||||||||||
OS_NTU_13 | 100.0/90.81 | 100.0/87.37 | 100.0/89.23 | 100.0/90.72 | 99.78/76.99 | 88.99/62.77 | 62.28/18.08 | 99.71/92.65 | 99.31/93.09 | |||||||||||
Snail_81R_02 | 95.62/79.79 | 95.03/76.55 | 92.73/71.82 | 96.15/88.76 | 92.08/59.95 | 72.37/30.19 | 36.09/6.03 | 98.67/89.93 | 99.50/73.61 | |||||||||||
Snail_81R_03 | 98.12/88.19 | 97.95/81.95 | 98.22/87.82 | 98.53/89.36 | 97.60/64.07 | 76.77/28.99 | 35.56/8.48 | 99.72/91.26 | 98.47/61.83 | |||||||||||
Garden_LT_01 | 98.40/87.89 | 87.30/79.01 | 87.68/82.31 | 75.35/64.21 | 74.40/63.33 | 48.96/39.26 | 86.10/39.30 | 94.76/81.20 | 76.18/35.97 | |||||||||||
Garden_LT_02 | 98.06/87.85 | 82.98/78.21 | 86.54/79.57 | 70.89/61.41 | 70.09/64.41 | 53.23/36.09 | 70.79/33.74 | 87.30/80.75 | 73.33/33.17 | |||||||||||
Garden_LT_03 | 95.08/83.31 | 94.44/79.91 | 90.19/80.34 | 84.94/77.60 | 67.55/46.29 | 53.70/43.79 | 61.88/28.28 | 89.20/81.52 | 78.73/47.56 | |||||||||||
Garden_LT_04 | 97.03/86.69 | 94.60/86.45 | 91.64/81.17 | 85.36/73.77 | 69.24/55.00 | 40.73/39.47 | 73.55/31.12 | 92.61/82.24 | 84.80/61.57 | |||||||||||